OpenStudy (anonymous):

Which of the following statements follows from (x - 3) 2 = 5? 1 = 2x 2 + 7x a.)40 B.)41 c.)57

jagr2713 (jagr2713):

\[ \huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }\] @golden00 are we suppose to solve for x?

OpenStudy (anonymous):

Thanks and i believe so not really sure @jagr2713

jagr2713 (jagr2713):

me to not sure but let me ask someone, @sleepyjess

OpenStudy (sleepyjess):

uuuhhh I am confused by this question :/

jagr2713 (jagr2713):

me to D:

OpenStudy (sleepyjess):

I am confused on what is supposed to be done

jagr2713 (jagr2713):

@golden00 can you take a screen shot of the question and send it to us?

OpenStudy (sleepyjess):

^^^^ That might help

OpenStudy (anonymous):

Does this help^

jagr2713 (jagr2713):

this is a test?

hartnn (hartnn):

yes it does :) Take square root on both sides let me give you an example : \(\Large a^2 = b \\ \Large a = \pm b\)

OpenStudy (anonymous):

so confused^